Given the national income identity Y=C+I+G+(X-M) and C=300+0.8Yd 1=650 G=750 X=550 M=350+0.15Y T=0.25Y
CALCULATE
(a) 1.injections=exports, investments
2.Withdrawals=imports and savings
(b) The marginal propensity to save on disposable income
(c) The marginal propensity to import is 0.15
(ci)The marginal propensity to tax is 0.25
(d)The marginal propensity to consume on national income
(e) The equilibrium national income level
(f) The change in equilibrium level if investment increases by 10%
(g)Change in government expenditure if full employment income level is 3600
1
Expert's answer
2019-10-22T09:37:26-0400

Y=C+I+G+(X-M) = 300+0.8 Y + 650+750+550-350-0.15Y=1900 + 0.65Y

0.35 Y = 1900

Y = 5428.5

M=350+0.15Y=1164.28

Net Export = 550 - 1164.28 = -614.28. this means that imports exceed exports by 614.28.

MPC = (300+0.8Y)/Y=0.85

MPS = 1-0.85 = 0.15

if investments increase by 10%, total income will increase by more than 10% due to the effect of the investment multiplier.

government spending will be reduced